Rauschenberg Manufacturing is investigating which locations would best position its new plant relative to three important customers​ (located in cities​ A, B, and​ C). As shown in the table​ below, all three customers require multiple daily deliveries. Management limited the search for this plant to those three locations and compiled the following​ information:

Location   Deliveries per day   x-coordinate   y-coordinate
A   3   300   300
B   9   300   100
C   2   200   200

a. Which of these three locations yields the smallest total travel​ distance, based on Euclidean​ distances?

Location yields the shortest travel​ (transportation) distance of nothing miles.
Answer #1

distance travelled when the plant is located at A  

3 x0 + 9x [ (300-300)2 + (300-100)2]1/2 + 2x[ (300-200)2+ ( 300-200)2]1/2= 1800+ 282.84 = 2082.84

Distance travelled when the plant is at B

3x [ (300-300)2+(100-300)2]1/2 +0+2 x[ (300-200)2+(100-200)2]1/2 = 600+282.84 = 882.84

Distance when plant is located at C

3 x{ (200-300)2 + (200-300)2]1/2+9x [ (300-200)2+(100-200)2]1/2+0 = 1697.04

The least cost is obtained when the location is at B. The distance is 882.84